How do you solve 2x/x-3 + x/x-8 = 3/x^2-11x+24?
Answer by stanbon(75887)

How do you solve 2x/x-3 + x/x-8 = 3/x^2-11x+24
-----------------------------------
Since (x-3)(x-8) = x^2-11x+24,
Multiply thru the equation by x^2-11x+24 to get:
----
2x(x-8) + x(x-3) = 3
2x^2-16x + x^2-3x = 3
3x^2 -19x -3 = 0
----
x = [19 +- sqrt(361 - 4*3*-3)]/6
---
x = [19 +- sqrt(388)]/6
---
x = [19 +- sqrt(4*97)]/6
---
x = [19 +- 2sqrt(97)]/6
================================
Cheers,
Stan H.